OpenStudy (anonymous):

The distance on the number line from -10 to -2 equals 5x – 12. What is the value of x?

OpenStudy (shane_b):

The total distance from -10 to -2 = 8 units. So \[5x-12=8\]Just solve for x.

OpenStudy (aliahmed):

x=2 or x=0.4

OpenStudy (anonymous):

Im checking on it..

OpenStudy (aliahmed):

why eq =8

OpenStudy (aliahmed):

??

OpenStudy (aliahmed):

(-10 or -2)=5x-12 yeah ??

OpenStudy (anonymous):

So if is -10-(-2)= -8 5x-12=-8 5x=4 4/5= -8 is that right???

OpenStudy (shane_b):

The distance is 8 because the question states: "distance on the number line from -10 to -2 *equals* 5x – 12" And the distance between -10 and-2 on a number line is just 8...just so that's clear.

OpenStudy (shane_b):

You should just end up with x=4.

OpenStudy (shane_b):

Here's it completely written out:\[5x-12=8\]\[5x=20\]\[x=4\]

OpenStudy (aliahmed):

-10 to -2 not meaning -10 -(-2) right?

OpenStudy (shane_b):



OpenStudy (aliahmed):

all eq meaning the value between -10 &-2

OpenStudy (shane_b):

The distance between the two points on the line will always be positive...and as you can see in the picture it's 8 units.

OpenStudy (aliahmed):

not satisfaction ^_^

OpenStudy (shane_b):

You could think of it this way: \[-2-(-10)=8\]

OpenStudy (aliahmed):

i believe it meaning interval value between the -10 to -2 such as -3 or -5 ,-8 & -9 ?? right

OpenStudy (shane_b):

There are an infinite number of values between -10 and -2 though...

OpenStudy (aliahmed):

i am sorry if discomfort you

OpenStudy (shane_b):

and it specifically says distance. You're not :)

OpenStudy (aliahmed):

yup i know that

OpenStudy (aliahmed):

u are right i know what u meaning my friend

OpenStudy (shane_b):

ok :)

OpenStudy (aliahmed):

yup yup ok

OpenStudy (aliahmed):

where u are from ??

OpenStudy (anonymous):

first i got x=4 and then he reply about negativity, thats where i got confused.. but anyway thank you.. and ok..